# Learning Primitive Embodied World Models: Towards Scalable Robotic Learning

### v2

@@ 4.2 Application 2: Plug-and-Play Long-Horizon Compositional Generalization
In the previous section, we described open-loop future prediction via single-step rollout. In this section, we close the loop by feeding the model’s generated future frames back as input for subsequent predictions, enabling iterative, on-the-fly planning and execution:
$\mathbf{x}_{t+1:T}^{(k)}\sim\mathcal{W}(\mathbf{x}_{t},h^{(k)}),\quad\mathbf{x}_{t}\leftarrow\text{crop}(\mathbf{x}_{t}^{(k)}),$
where $k$ denotes the current planning step, and the cropped latest frame from the generated rollout becomes the new input $\mathbf{x}_{t}$ for the next prediction. This autoregressive, closed-loop deployment allows the world model to continuously adapt to the actual environment state, correcting for drift and disturbances. Figure 5 presents the pipeline of this hierachical long-horizon compositional generalization. Appendix C presents a more detailed workflow.
@@
Table 1: Overall success rate on RLBench tasks. We compare our method against existing baselines on 9 manipulation tasks from the RLBench benchmark (James et al., 2020b), with success rates averaged over 100 episodes. Results for Image-BC (Jang et al., 2022), UniPi (Du et al., 2023), and 4DWM (Zhen et al., 2025a) are directly cited from the 4DWM paper. Our method (bottom row) achieves the highest success rate on most tasks, with consistent gains in articulated and contact-rich scenarios.
@@ 4.3 Application 3: As Data Synthesis Engines
Beyond direct policy execution, our Primitive World Model (PWM) serves as a powerful data synthesis engine for scalable embodied learning. By conditioning on diverse language commands and initial visual states, the model can generate vast amounts of physically plausible, high-fidelity video rollouts that exhibit consistent robot dynamics. This capability is particularly valuable for addressing the critical data bottleneck in robotics. We can synthesize data for rare or hard-to-collect scenarios (e.g., specific object interactions, failure modes) or augment existing datasets with variations in object appearance, lighting, and background, thereby enriching the training corpus for downstream tasks like imitation learning or reinforcement learning. Crucially, because our model is trained on a mixture of real and simulated data and learns soft embodiment-aware constraints from full-arm observations, the generated videos possess a high degree of physical consistency (EPiCS score: 11.45/13, see Table 12), making them more suitable for sim-to-real transfer than purely graphics-based simulation. This application transforms the world model from a planning tool into a force multiplier for data-driven robotics research.
4.4 Other Promising Applications
The modular and explainable nature of our framework unlocks several other promising applications. First, it enables Latent Action Prediction, where the generated video rollouts can be distilled into compact, interpretable latent action codes that capture the essence of a manipulation primitive, facilitating efficient communication and storage. Second, the framework can be extended towards Unified Generation of Action, Reward, or Camera Parameters. For instance, by introducing auxiliary heads or conditioning mechanisms, the same underlying model could potentially predict not only future frames but also estimate the expected reward for a generated trajectory or suggest optimal camera viewpoints for better scene understanding, creating a more holistic planning system. Finally, the generated videos provide a natural medium for Human-in-the-Loop Interaction and AR-based Planning. A human operator could view the predicted video sequence in an augmented reality (AR) interface, provide real-time feedback to correct or refine the plan before execution, or even interactively edit the video to specify desired outcomes, creating a more intuitive and collaborative robotics interface. These applications highlight the framework’s potential as a foundational component for next-generation, human-centered robotic systems.
@@ 5 Analysis
5.1 Performance
Planning Our method performs a notable success rate, attributable to the effective decomposition of tasks into manageable primitives.
The test scenarios outlined in Table 2 were not present in the training set, yet our method achieves a high primitive accuracy of 18 out of 20 for the task of picking up a cup. Similar performance is observed across other tasks, with accuracy rates of 16 out of 20 for moving cloth and 15 out of 20 for folding cloth. This robust planning capability ensures that the robot can generate reliable sequences of actions, setting a strong foundation for successful execution. Notably, OpenVLA fails completely in the zero-shot setting, highlighting its reliance on task-specific adaptation. In contrast, our method achieves strong generalization without retraining the video model, benefiting from modularity and spatial grounding in its design.
@@ 5.1 Performance
Primitive Execution The results further highlight the effectiveness of our approach. The integration of planning and video generation significantly contributes to the overall task success rates. For instance, in the task of picking up a cup, our method achieves a task success rate of 16 out of 20, compared to only 12 out of 20 for the OpenVLA baseline. Similar trends are observed in the other tasks, with our method outperforming OpenVLA in both moving and folding cloth. Notably, the ability to succeed in these tasks–despite their absence in the training set–demonstrates the robustness and generalization capability of our method. This enhanced execution performance can be attributed to the careful segmentation of primitives and the use of high-quality video generation, which collectively facilitate precise and effective task completion.
Long-Horizon Tasks The proposed methodology effectively addresses long-horizon tasks by decomposing them into a series of modular primitives, each designed to be plug-and-play. This hierarchical framework enables each primitive to concentrate on short-range objectives, which aligns closely with the strengths of Vision-Language Models (VLM). By constraining each component to operate within a limited contextual scope, we mitigate potential challenges such as contextual inconsistencies that may arise during long-term planning. As illustrated in Appendix I.4 Figure 13, the process of mapping generated video frames to real-world coordinates through the camera’s intrinsic and extrinsic parameters underpins this modular approach. This mapping facilitates precise pose estimations for each primitive, thereby ensuring that the robotic system can execute tasks effectively without losing sight of the overarching long-horizon goal. Consequently, our methodology leverages the inherent advantages of decomposing complex tasks into manageable segments, enhancing both efficiency and reliability in execution.
5.2 Efficiency
As shown in Appendix Table 8, our model demonstrates significant advantages in both computational efficiency and memory footprint. Compared to large-scale diffusion baselines such as Hunyuan I2V and Wan 2.1, our method achieves up to 75x faster inference and 6–7x lower VRAM usage, making it well suited for real-time robotic deployment. Despite its lightweight design, our method maintains competitive video quality and task relevance, validating its practicality in resource-constrained scenarios.
5.3 Ablation Study
@@
Ablation Group Variant Pick up cup Move cloth Fold cloth
Primitive Planner Full model (with start/end prompts) 16 / 20 14 / 20 13 / 20
w/o start-end prompt 12 / 20 10 / 20 7 / 20
w/o primitive planner (direct instruction-to-action) 9 / 20 5 / 20 3 / 20
Video Generation Full model (with sim + real data) 16 / 20 14 / 20 13 / 20
Trained on real-only data 12 / 20 9 / 20 5 / 20
We conduct two ablation studies to assess the contribution of (1) the VLM-based primitive planner, and (2) simulation-augmented training for video generation. As shown in Table 3, removing the start-end spatial prompts from the planner leads to a noticeable decline in performance across all tasks. This confirms that explicitly grounding primitives with spatial anchors significantly improves downstream execution. A further removal of the entire primitive decomposition step–i.e., mapping instructions directly to actions–results in severe performance degradation, especially on multi-step or spatially entangled tasks like cloth folding. For video generation, removing simulated data from the training corpus substantially weakens the model’s rollout quality, leading to lower execution success rates. This suggests that simulation data offers critical diversity and coverage, allowing the model to generalize better to real-world scenes with varied configurations and object appearances.
@@ 6 Related Work
Video Generation as World Models Given the overwhelming success of Diffusion Models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021; 2022) in image generation (Rombach et al., 2022b; Ramesh et al., 2022; Ding et al., 2022), video diffusion models (Li et al., 2022; Ho et al., 2022; He et al., 2023; Li et al., 2023a; Blattmann et al., 2023; Singer et al., 2022; Wang et al., 2023; Chen et al., 2023; 2024a; Xing et al., 2023) have emerged as a prominent research focus. With the emergence of Sora (Brooks et al., 2024; Liu et al., 2024e), Transformer-based video generation has gained traction as a mainstream paradigm. Leveraging the Diffusion Transformer (DiT) architecture Peebles & Xie (2023), these models significantly improve video stability and temporal consistency (Zheng et al., 2024; Lin et al., 2024; Yang et al., 2024c; Kong et al., 2024; Polyak et al., 2025; Yang et al., 2025b; Wang et al., 2025). Empowered by strong visual priors, video diffusion models show promising potential as pixel-level world models. A world model learns to predict future system states, either in a low-dimensional latent space (Achille & Soatto, 2017; Ferns et al., 2004; Lesort et al., 2018) or directly in pixel space (Chiappa et al., 2017; Ha & Schmidhuber, 2018; Hafner et al., 2020; Micheli et al., 2022; Chen et al., 2022). A large body of work (Ko et al., 2023; Du et al., 2023; Yang et al., 2023; Brooks et al., 2024; Black et al., 2024a; Xiang et al., 2024; Bruce et al., 2024; Hu et al., 2024; Zhou et al., 2024; Dalal et al., 2025; Liang et al., 2024a; Clark et al., 2025; Hu et al., 2024; Li et al., 2025; Team et al., 2025a; Pertsch et al., 2025; Zhen et al., 2025a; Zhu et al., 2025; Hafner et al., 2025) has framed video generation as a new methodology for learning world models. Latest works, typically DiT-based and following a unified generation paradigm (Li et al., 2025; Team et al., 2025a; Zhen et al., 2025a), are costly to train and lack real-time responsiveness. While hierarchical approaches (Ajay et al., 2023; Dalal et al., 2025; Zhou et al., 2024) offer structure, they struggle with real-time adaptation in dynamic environments and rigid cross-modal alignment. Error accumulation in layered designs remains an open problem. Several approaches (Ajay et al., 2023; Dalal et al., 2025; Zhou et al., 2024) adopt hierarchical architectures, among which Ajay et al. (2023); Dalal et al. (2025) are the closest to ours. However, these methods lack the integration of vision-language reasoning, planning, and grounding as in our approach. In contrast, our approach is distinguished by its ability to perform zero-shot policy execution, without relying on task-specific trained policies that often suffer from limited generalizability. Furthermore, our start-to-end point guidance surpasses their interactive refinement or reliance on 3D models in both efficiency and explainability. Some additional studies Seo et al. (2023); Wu et al. (2024); Yang et al. (2024b); Cheang et al. (2024) investigated non-diffusion approaches to modeling embodied world models.
End-to-End Vision-Language-Action Learning End-to-end Vision-Language-Action (VLA) models constitute another major branch of embodied learning, directly mapping visual and linguistic inputs to actions without explicit dynamics modeling. Recent advances in vision-language models (VLMs) Chen et al. (2024c); Li et al. (2024a); Beyer et al. (2024); Li et al. (2024b) have catalyzed the development of unified VLA systems that integrate perception, grounding, reasoning, and control within a single architecture–departing from traditional modular pipelines. At the core of most VLA systems lies the intuition that pre-trained vision-language models (VLMs) can serve as high-level semantic engines. Methods like RT-2 Brohan et al. (2023b), OpenVLA Kim et al. (2024a), and RoboAgent Bharadhwaj et al. (2023) inject Internet-scale knowledge into robot policies by fine-tuning VLMs or conditioning them with low-level control tokens. Based on their action modeling paradigms, VLA approaches can be broadly categorized into regression-based Li et al. (2023b); Brohan et al. (2023a); Liu et al. (2024c), generative/diffusion-based Black et al. (2024b); Liu et al. (2024d); Team et al. (2024), and hybrid methods Liu et al. (2025a); Ye et al. (2024b). While these unified policies show strong performance on short-horizon or single-task scenarios, they often struggle with generalization and scalability–particularly across long horizons, domains, or embodiments. To address these limitations, recent efforts have introduced hierarchical structures into end-to-end VLA models, either by modularizing perception and policy Liu et al. (2024b); Team et al. (2025b); Bjorck et al. (2025), or by structuring execution into multi-stage skills Shi et al. (2025); Physical Intelligence (2025); Raj et al. (2024). Though often labeled hierarchical, many still preserve end-to-end differentiability or latent consistency across stages. Such designs improve sample efficiency, handle long-horizon dependencies, and provide interpretable intermediate reasoning. Despite these advances, existing VLA systems remain heavily dependent on in-domain demonstrations. Some mitigate this via cross-modal or cross-embodiment data Intelligence et al. (2025); NVIDIA et al. (2025), or by scaling human video demonstrations Yang et al. (2024a); Tian et al. (2024), yet diversity and coverage remain limited. Our work enables zero-shot generalization across unseen tasks and settings, while maintaining modularity, interpretability, and execution efficiency.
@@
Datasets for Robotic Learning High-quality and diverse data is essential for generalizable robotic learning Xue et al. (2025b). Existing large-scale datasets are typically collected via human teleoperation Brohan et al. (2022); Dass et al. (2023); Khazatsky et al. (2024a), scripted pipelines Collaboration et al. (2023); Gu et al. (2023), or expert policies Schiavi et al. (2023), and have largely been consolidated in OpenX-Embodiment Collaboration et al. (2023). Despite its scale, this benchmark remains costly to scale and limited in semantic and morphological diversity. To enhance compositionality, RH20T-p Chen et al. (2024b) introduced primitive-level annotations atop RT-20T Fang et al. (2023), but suffers from label noise and redundancy. Meanwhile, simulated benchmarks such as RLBench James et al. (2020b), CALVIN Mees et al. (2022), and LIBERO Liu et al. (2023) offer controllable environments, though often lack photorealism or task variety. Recent works Mandlekar et al. (2023); Garrett et al. (2024); Yang et al. (2025a); Xue et al. (2025a) explore augmenting human demos via simulation or generative rendering, but still rely on expert trajectories. In contrast, our framework directly generates large-scale, primitive-level demonstrations in simulation without human supervision or segmentation. Combined with pre-trained vision-language models, this enables efficient sim-to-real transfer with strong generalization.
@@ 7 Conclusion, Limitations, and Further Vision
We present a modular and interpretable framework for zero-shot embodied policy learning, combining a vision-language planner with a primitive-conditioned video diffusion world model. By grounding high-level instructions into spatially guided primitives and reusing a lightweight video model across tasks, our system achieves strong performance on challenging real-world manipulation benchmarks. Extensive experiments demonstrate superior generalization, data efficiency, and runtime scalability. Our work highlights the potential of combining semantic reasoning with visual dynamics modeling for scalable, reusable, and task-agnostic robotic intelligence.
7.1 Brief Summary of Methodology Advantages
Against Vanilla Video Diffusion Models-Based Methods Compared to long-horizon video diffusion models, our method operates at the primitive level with short-horizon prediction, drastically reducing computational cost and optimization difficulty. Leveraging causal distillation and self-forcing (Huang et al., 2025), we achieve real-time generation (12 FPS) in just 4 denoising steps, enabling online, on-the-fly rollout. This causal, modular design avoids error accumulation and supports dynamic, adaptive control–critical for real-world robotic deployment.
Against VLA Models Unlike end-to-end Vision-Language-Action (VLA) models that act as black boxes, our framework is modular and interpretable. By decoupling high-level planning (via a VLM) from low-level dynamics (via a diffusion world model), we enable transparent, debuggable execution. This design supports zero-shot generalization without task-specific fine-tuning–whereas VLA models often fail in such settings–and allows flexible integration of symbolic control (e.g., gripper actions), enhancing robustness and safety.
Against Imitation Learning In contrast to imitation learning, which relies on costly, task-specific demonstration data, our method enables compositional generalization through primitive-based learning. By decomposing tasks into reusable, semantically meaningful units, we support both intra-primitive (concept-level) and inter-primitive (sequence-level) generalization. Moreover, our hybrid sim-real data strategy allows leveraging unlabeled video data at scale, breaking the dependency on paired state-action data and paving the way for web-scale embodied learning.
7.2 Limitations and Further Vision
First, the current pipeline relies on a VLM for high-level planning, forming a semi-closed loop. A natural evolution is to develop a unified, real-time understanding-and-generation system, where action generation and perception are seamlessly integrated. This could enable richer online interaction–such as real-time user feedback during autoregressive (AR) video generation–and potentially allow planning directly in the model’s latent space, bypassing explicit language instructions altogether.
Second, while our 4-step causal distillation achieves real-time performance (12 FPS), further latency reduction is needed for high-frequency control. Emerging AR video models offer a promising path toward better speed-quality trade-offs. Additionally, our evaluation is currently limited to single-arm rigid-body manipulation. Extending the approach to dual-arm coordination, multi-agent systems, or deformable object manipulation (e.g., cloth folding) remains an open challenge.
Finally, despite the gains in data efficiency, a standardized, scalable benchmark for embodied world models is still lacking. We view our primitive-centric, hybrid sim-real framework not just as a model, but as a foundational paradigm, yet increasingly vital for building general, interpretable, and scalable embodied intelligence.
@@ 7.2 Limitations and Further Vision
Appendix A Theoretical Justification for the Compositional Expressivity of Motion Primitives
We present a formal, mathematically rigorous derivation of the expressivity of finite motion primitives. Rather than treating compositionality as an empirical observation, we derive it from three minimal, observationally grounded axioms using tools from topology, dynamical systems, and computability theory. The resulting framework is logically consistent, internally complete, and resistant to counterarguments due to the simplicity and universality of its assumptions.
A.1 Foundational Axioms: Irrefutable Empirical Premises
We begin with three axioms, each grounded in direct empirical observation and robust across biological and artificial agents.
Axiom 1 (Finiteness of Actuation).
The agent’s low-level action space is finite and discrete: $\mathcal{P}$ is a finite set.
Justification: All physical agents have finite degrees of freedom, finite control bandwidth, and finite resolution in actuation (e.g., motors, grippers). Even continuous control policies are ultimately discretized at the hardware level. This is a consequence of the finite energy and material constraints of embodied systems.
Axiom 2 (Continuity of Embodiment).
Embodied trajectories evolve continuously in a compact metric space $(X,d)$, where $X\subset\mathbb{R}^{n}$ represents the configuration or sensory manifold. That is, any behavior $\xi\in\mathcal{D}_{\text{emb}}$ is a continuous path:
$\xi:[0,T]\to X,\quad\text{with }T>0.$
Justification: Physical motion is governed by differential equations (Newtonian mechanics, motor control), hence trajectories are continuous (Lipschitz or smoother). No real robot or animal exhibits teleportation or discontinuous state jumps.
Axiom 3 (Observational Fidelity).
For any two distinct behaviors $\xi_{1}\neq\xi_{2}$ in $\mathcal{D}_{\text{emb}}$, there exists a finite sequence of observations that distinguishes them with arbitrary precision.
Justification: Perception systems (cameras, proprioception) can resolve differences in trajectories up to sensor resolution. This ensures that $\mathcal{D}_{\text{emb}}$ is separable and observable.
A.2 Formal Framework: Topological Dynamical Systems
Let $\mathcal{P}=\{f_{1},\dots,f_{n}\}$ be a finite set of partial continuous maps $f_{k}:X\rightrightarrows X$, each representing the state transition induced by executing primitive $a_{k}$. Assume each $f_{k}$ has compact support and finite duration.
Let $\mathcal{M}(X)$ denote the space of continuous paths in $X$ over finite time intervals, equipped with the compact-open topology (uniform convergence on compact sets).
Define the orbit semigroup generated by $\mathcal{P}$ as:
$\langle\mathcal{P}\rangle_{+}=\left\{f_{i_{m}}\circ\cdots\circ f_{i_{1}}\mid m\in\mathbb{N},\ i_{j}\in\{1,\dots,n\}\right\},$
the set of all finite compositions of primitive transitions.
Each composition induces a trajectory $\xi\in\mathcal{M}(X)$ via concatenation:
$\xi(t)=\begin{cases}f_{i_{1}}(x_{0},t),&t\in[0,\tau_{1})\\ f_{i_{2}}(\xi(\tau_{1}),t-\tau_{1}),&t\in[\tau_{1},\tau_{1}+\tau_{2})\\$ $\vdots\end{cases}$
Let $\mathcal{C}(\mathcal{P})\subset\mathcal{M}(X)$ denote the set of all such composite trajectories.
A.3 Key Lemma: Density via Topological Transitivity
Lemma A.1 (Density of Composite Orbits).
Suppose the action of $\langle\mathcal{P}\rangle_{+}$ on $X$ is topologically transitive and $X$ is a compact connected manifold. Then $\mathcal{C}(\mathcal{P})$ is dense in $\mathcal{M}(X)$ under the compact-open topology.

### v3

@@ 4.2 Application 2: Plug-and-Play Long-Horizon Compositional Generalization
In the previous section, we described open-loop future prediction via single-step rollout. In this section, we close the loop by feeding the model’s generated future frames back as input for subsequent predictions, enabling iterative planning and execution: $\mathbf{x}_{t+1:T}^{(k)}\sim\mathcal{W}(\mathbf{x}_{t},h^{(k)}),\quad\mathbf{x}_{t}\leftarrow\text{crop}(\mathbf{x}_{t}^{(k)})$, where $k$ denotes the current planning step, and the cropped latest frame from the generated rollout becomes the new input $\mathbf{x}_{t}$ for the next prediction. This autoregressive, closed-loop deployment allows the world model to continuously adapt to the actual environment state, correcting for drift and disturbances. Figure 5 presents the pipeline of this hierachical long-horizon compositional generalization. Appendix C presents a more detailed workflow.
@@
Table 1: Overall success rate on RLBench tasks. We compare our method against existing baselines on 9 manipulation tasks from the RLBench benchmark (james2019rlbench), with success rates averaged over 100 episodes. Results for Image-BC (jang2022bc), UniPi (du2023learning), and 4DWM (zhen2025tesseract) are directly cited from the 4DWM paper. Our method (bottom row) achieves the highest success rate on most tasks, with consistent gains in articulated and contact-rich scenarios.
@@ 4.3 Application 3: As Data Synthesis Engines
Beyond policy execution, our Primitive World Model serves as a scalable data synthesis engine. By conditioning on language instructions and initial visual states, it generates physically plausible, high-fidelity video rollouts with consistent robot dynamics–enabling data augmentation for rare interactions, failure modes, or scene variations (e.g., lighting, background). Trained on hybrid sim-real data and full-arm observations, the model internalizes soft embodiment constraints, yielding videos with strong physical consistency (EPiCS: 11.45/13, Table 12). This makes the generated data more effective for sim-to-real transfer than graphics-based simulation, turning the world model into a force multiplier for imitation and reinforcement learning.
For more promising applications, please see Section 7 and Appendix  L.
@@ 5 Analysis
Performance1) Planning. Our method achieves high planning accuracy on unseen tasks by decomposing them into spatially grounded primitives. As shown in Table 2, it attains 18/20 primitive accuracy on “pick up cup,” 16/20 on “move cloth,” and 15/20 on “fold cloth.” In contrast, OpenVLA fails completely in zero-shot (0/20), underscoring its dependence on task-specific fine-tuning. Our approach generalizes without retraining the video model, thanks to its modular design and explicit spatial grounding. 2) Primitive Execution. The integration of planning and video generation translates into strong execution: our method achieves 16/20 task success on cup picking, versus 12/20 for OpenVLA; similarly, 14/20 vs. 10/20 for moving cloth, and 13/20 vs. 4/20 for folding cloth. Success on these out-of-distribution tasks demonstrates robust generalization, enabled by precise primitive segmentation and high-fidelity video rollouts. 3) Long-Horizon Tasks. We handle long-horizon tasks by decomposing them into short-horizon, plug-and-play primitives—each aligned with the contextual scope of vision-language models (VLMs). This mitigates drift and inconsistency in extended planning. As shown in Appendix I.4 (Fig. 13), generated video frames are projected to real-world coordinates via camera intrinsics/extrinsics, enabling accurate pose estimation per primitive while preserving global task coherence.
@@
Primitive-level Compositional Generalization  A core strength of our primitive-centric framework is its ability to achieve conceptual compositional generalization (Figure 3)–the capacity to succeed on novel tasks that were never explicitly demonstrated during training. We quantitatively evaluate this capability in Table 3, which reports success rates for executing unseen (predicate, object) combinations on a real robot. The results demonstrate strong generalization: for instance, the "pick" primitive, trained on objects like cups and boxes, successfully generalizes to the novel combination "pick jar" with an 80% success rate. Similarly, the "push" action generalizes effectively to cups and jars, despite these pairings being absent from the training data. Even for the more complex "open" action, the model transfers successfully to the unseen "open cup" task (70% success), showcasing its ability to adapt a skill to an object with significantly different geometry and affordances. This generalization is enabled by two key factors. (1) Our video diffusion model, trained on densely annotated primitive data, learns rich, disentangled representations of action dynamics and object properties. This allows it to condition the generation of a "pick" motion on the visual features of a previously unseen jar. (2) The VLM-based primitive planner provides semantic and spatial grounding, correctly identifying the target object and generating a feasible subgoal configuration even for unseen objects. The high success rates on these zero-shot combinations underscore that our approach moves beyond simple imitation learning, instead capturing the underlying compositional structure of embodied manipulation, a crucial step toward truly flexible and scalable robotic intelligence.
Predicate \ Object cup box drawer jar
pick 9/10 8/10 8/10
open 9/10 9/10 7/10
push 7/10 8/10 8/10 6/10
Efficiency  Our method is significantly more efficient than large diffusion baselines (e.g., Hunyuan I2V, Wan 2.1), achieving up to 75x faster inference and 6–7x lower VRAM usage (Appendix Table 9). This enables real-time deployment without sacrificing video quality or task relevance.
Ablation Study   We ablate two core components: (1) the VLM-based primitive planner, and (2) simulation-augmented video training. Results in Table 4 show that removing Start-Goal Guidance reduces success rates (e.g., from 16/20 to 12/20 on cup picking), confirming the value of explicit spatial grounding. (3) Eliminating primitive decomposition entirely—mapping instructions directly to actions—causes severe drops (e.g., 3/20 on cloth folding), highlighting the necessity of structured planning. Similarly, training the video model on real data alone (without simulation) degrades performance across all tasks, verifying that simulated data enhances generalization to real-world object configurations.
@@
Ablation Group Pick up cup Move cloth Fold cloth
Full model 16 / 20 14 / 20 13 / 20
w/o SGG 12 / 20 10 / 20 7 / 20
w/o Primitive Planner 9 / 20 5 / 20 3 / 20
w/o Sim data 12 / 20 9 / 20 5 / 20
@@ 6 Related Work
Video Generation as World Models. Diffusion models have become dominant in image generation (ho2020denoisingdiffusionprobabilisticmodels; song2021scorebasedgenerativemodelingstochastic; rombach2022highresolutionimagesynthesislatent), and their extension to video—via 3D convolutions, latent spaces, or spatiotemporal attention—has yielded strong generative performance (he2023latentvideodiffusionmodels; blattmann2023stablevideodiffusionscaling; singer2022makeavideotexttovideogenerationtextvideo; chen2024videocrafter2). The success of Sora (videoworldsimulators2024; liu2024sora) has further established Transformer-based video diffusion (e.g., DiT (peebles2023scalablediffusionmodelstransformers)) as a leading paradigm, enabling improved temporal coherence and scalability (zheng2024opensorademocratizingefficientvideo; yang2024cogvideox; kong2024hunyuanvideo). These models are increasingly viewed as pixel-level world models—systems that predict future observations given actions and context (ha2018recurrent; hafner2020mastering; micheli2022transformers). Recent works frame video generation itself as world modeling (Ko2023Learning; du2023learning; bruce2024genie; zhou2024robodreamerlearningcompositionalworld; li2025unified), often using large DiT-based architectures trained under unified generative objectives. However, such models are computationally expensive and lack real-time responsiveness. Hierarchical alternatives (ajay2023compositional; dalal2024manipgen; zhou2024robodreamerlearningcompositionalworld) improve structure but suffer from error propagation and rigid cross-modal alignment. Crucially, none integrate vision-language reasoning, planning, and grounding in a zero-shot policy execution framework—key distinctions of our approach.
End-to-End Vision-Language-Action Learning. Vision-Language-Action (VLA) models bypass explicit dynamics modeling by directly mapping multimodal inputs to actions (rt-2; kim2024openvla; bharadhwaj2023roboagent). Leveraging pre-trained vision-language models (VLMs) (chen2024internvl; li2024llavaov; beyer2024paligemma), they unify perception, grounding, and control in a single architecture. Action prediction is typically implemented via regression (brohan2023rt; liu2024robomamba), diffusion (black2024pi_0; liu2024rdt), or hybrid schemes (liu2025hybridvla; ye2024latent). While effective on short-horizon tasks, these systems struggle with long-horizon generalization and cross-domain transfer. Recent efforts introduce hierarchy—via modular perception-policy stacks (liu2024self; bjorck2025gr00t) or skill decomposition (shi2025hi; raj2024learning)—to improve scalability and interpretability. Yet, nearly all rely on in-domain demonstrations, limiting zero-shot adaptability. Our method circumvents this dependency by generating task-agnostic, primitive-level trajectories in simulation, enabling zero-shot policy execution without task-specific fine-tuning.
@@
Datasets for Robotic Learning. Large-scale robotic datasets—collected via teleoperation (rt-1; khazatsky2024droid), scripted policies (open_x_embodiment_rt_x_2023; gu2023maniskill2), or expert rollouts (schiavi2023learning)—have been consolidated in benchmarks like OpenX-Embodiment (open_x_embodiment_rt_x_2023). Despite scale, they remain costly to expand and limited in semantic or morphological diversity. RH20T-p (chen2024rh20tp) adds primitive annotations but introduces noise; simulated environments (e.g., RLBench (james2019rlbench), CALVIN (mees2022calvin)) offer control but lack realism. Recent works augment data via simulation or generative rendering (mimicgen; robosplat; demogen), yet still depend on human demonstrations. In contrast, our framework synthesizes diverse, primitive-level trajectories de novo in simulation—without human supervision or segmentation—enabling efficient sim-to-real transfer when paired with pre-trained VLMs.
@@ 7 Conclusion, Limitations, and Further Vision
We propose a novel perspective on embodied world modeling: using unlabeled action videos as a universal representation, we segment long-horizon trajectories into semantically indivisible primitives–short clips aligned with atomic language instructions. This shift unlocks three key advantages: (1) Fine-grained language–vision alignment enables high-fidelity, efficient video generation with lower compute; (2) Primitive-centric data is lower-dimensional, denser, and easier to collect, while diffusion models provide strong intra-primitive compositional generalization; (3) A VLM planner can sequence these primitives via Start-Goal Heatmap Guidance (SGG), achieving inter-primitive “building-block” generalization–turning the world model into both a data engine and a plug-and-play policy backbone.
Our method outperforms imitation learning, vanilla video diffusion, VLAs, and state-of-the-art embodied world models by combining modular, primitive-level modeling with self-forcing distillation for real-time, low-step generation (12 FPS in 4 steps), zero-shot compositional generalization, and sim-to-real scalability–all while remaining interpretable, adaptive, and free from costly demonstrations or black-box policies, making it ideal for real-world robotic deployment.
While our framework marks a leap in real-time, interpretable robotic control, key frontiers remain: evolving toward a unified perception-action system that plans in latent space without language, pushing latency below 4-step inference for high-frequency tasks, and scaling to multi-agent, deformable-object, and dual-arm scenarios. Beyond model improvements, we call for standardized benchmarks to advance the field–positioning our primitive-centric, sim-real hybrid approach as a foundational paradigm for scalable, general-purpose embodied intelligence.
@@
Appendix A Theoretical Justification for Primitive-Centric Embodied Learning
We provide a mathematically grounded justification for the feasibility and expressivity of modeling embodied intelligence via a finite set of semantic primitives. Our derivation avoids pathological assumptions about trajectory indecomposability and instead builds upon three empirically plausible premises.
A.1 Foundational Assumptions
Let $X\subset\mathbb{R}^{d}$ be a compact metric space representing the robot’s observable state (e.g., RGB-D image, proprioception). Let $\mathcal{U}$ be a countable set of natural language instructions describing manipulation intents (e.g., “pick cup”, “arrange flowers”).
We assume:
Assumption A.1 (Semantic Grounding).
There exists a surjective mapping $\Phi:\mathcal{U}\to\mathcal{T}$, where $\mathcal{T}=\{\tau_{1},\dots,\tau_{K}\}$ is a finite set of semantic primitive templates, such that each $\tau_{k}$ corresponds to a reusable skill (e.g., “grasp”, “place”, “rotate”) that generalizes across objects and scenes.
Assumption A.2 (Temporal Locality).
Each execution of a primitive template $\tau_{k}$ under instruction $u=\Phi^{-1}(\tau_{k})$ yields a trajectory $p\in C([0,T_{k}],X)$ with bounded duration $T_{k}\leq T_{\max}$ (e.g., 2 seconds), and the mapping from initial state $x_{0}$ and goal $g$ to $p$ is continuous.
Assumption A.3 (Compositional Execution).
Any long-horizon task $U\in\mathcal{U}^{*}$ (finite sequence of instructions) induces a trajectory $\xi\in C([0,T],X)$ that can be written as a concatenation:
$\xi=p_{1}\cdot p_{2}\cdot\dots\cdot p_{N},$
where each $p_{i}$ is an instance of some $\tau_{k_{i}}\in\mathcal{T}$, grounded via perception.
A.2 Constructive Approximation Theorem
Let $\mathcal{M}_{T}(X)=\{\xi\in C([0,T],X)\mid T>0\}$ be the space of all finite-horizon continuous trajectories, equipped with the uniform metric:
$d_{\infty}(\xi_{1},\xi_{2})=\sup_{t\in[0,T]}\|\xi_{1}(t)-\xi_{2}(t)\|.$
Let $\mathcal{C}(\mathcal{T})\subset\mathcal{M}(X)$ denote the set of all trajectories obtainable by finite composition of primitive instances from $\mathcal{T}$.
Theorem A.4 (Density of Primitive Compositions).
Under Assumptions A.1A.3, for any $\xi^{*}\in\mathcal{M}_{T}(X)$ and any $\epsilon>0$, there exists a composed trajectory $\xi\in\mathcal{C}(\mathcal{T})$ such that $d_{\infty}(\xi,\xi^{*})<\epsilon$.